OpenStudy (anonymous):

The sum of two numbers is 20. Their difference is 4. a. Write a system of equations that describes this situation. b. Solve by elimination to find the two numbers

OpenStudy (campbell_st):

x + y = 20 x - y = 4

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

and idk how to solve it if the x and y values are the same in both equations @campbell_st

OpenStudy (campbell_st):

let x be the larger number, y the smaller number then there sum is 20 x + y = 20 their difference is 4 x - y = 4 to solve add the equations together... one letter will disappear and you have an equation with one unknown

OpenStudy (anonymous):

someone help me!

OpenStudy (anonymous):

I KEEP GETTING THE WRONG ANSWER:(

OpenStudy (anonymous):

PLEASE HELP1

OpenStudy (anonymous):

@helpmeplz310

OpenStudy (anonymous):

yes, xokatexo?

OpenStudy (anonymous):

please help!!!!!!1

OpenStudy (anonymous):

sorry, idk anything about that